# Stability of the Kim–Milman flow map

### v1

@@ 3.2.3 Distributions with asymptotically positive convex profiles
As a final example, we turn to a family of distributions introduced in [8, 7]. To define said family, we require the following definitions. For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we define the integrated convexity profile of $f$, denoted $\kappa_{f}:\mathbb{R}_{+}\to\mathbb{R}\cup\{-\infty\}$, to be
$\kappa_{f}(r)\coloneqq\inf\Bigl\{\frac{\langle\nabla U(x)-\nabla U(y),x-y\rangle}{\|x-y\|^{2}}\ :\ \|x-y\|=r\Bigr\}\,.$
We introduce this definition with the following example.
@@
If $R_{V}=0$, then $\kappa_{V}(r)\geq\alpha_{V}>0$ is equivalent to strong convexity over all of $\mathbb{R}^{d}$. Otherwise for $R_{V}>0$, the potential has integrated convexity profile which might be negative inside $B(0,R_{V})$, and while remaining strongly convex outside the ball.
@@
    the right-derivative of $\hat{g}$ at the origin exists, denoted $\hat{g}^{\prime}(0)$.
This function class leads to our final corollary.
@@
Let ${\alpha}>-1$ and suppose $\hat{g}\in\mathcal{G}$. Suppose $\mu\propto\gamma\exp(-h)$ where the integrated convexity profile of $h$ satisfies $\kappa_{h}(r)\geq\alpha-r^{-1}\hat{g}(r)$. Then,
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\frac{1}{1+\alpha}\exp\Bigl(\frac{3\hat{g}^{\prime}(0)}{2\,(1+\alpha)}\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
@@
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
$\qquad\qquad=\int_{\exp(-T)}^{1}(1+\alpha-\alpha r^{2})^{-3/2}\exp\Bigl(-\frac{3\hat{g}^{\prime}(0)}{2\alpha}\bigl((1+\alpha-\alpha r^{2})^{-1}-1)\bigr)\Bigr)\,\mathrm{d}r$
$\qquad\qquad\leq\exp\Bigl(-\frac{3\hat{g}^{\prime}(0)}{2\alpha}\bigl((1+\alpha)^{-1}-1)\bigr)\Bigr)\int_{\exp(-T)}^{1}(1+\alpha-\alpha r^{2})^{-3/2}\,\mathrm{d}r\,$,
@@
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\alpha^{-1}\exp\bigl(O\bigl((L_{V}/\alpha)\,(1\vee L_{V}R_{V}^{2})\bigr)\bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
@@ 4 Extension to stronger metrics
To start, we follow the start of the proof of Proposition 3.1 assuming ($\Theta$). It is easy to see that $\gamma$-almost surely, it holds that
@@
Carrying out the rest of the argument as before, we obtain $\gamma$-a.s.
$\|X_{T}-Y_{T}\|\leq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,\eta_{T}\coloneqq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\int_{0}^{T}d_{T-s}\exp\Bigl(\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$ (4.3)
@@
To this end, for $y\in\mathbb{R}^{d}$ and $t\geq 0$, we write
@@
${\rm KL}(\nu\,\|\,\mu)\coloneqq\int\log\frac{\mu}{\nu}\,\mathrm{d}\mu\leq\frac{\lambda}{2}\,{\rm FI}(\nu\,\|\,\mu)\,.$
@@
$\nabla\log\frac{\nu Q_{s}}{\mu Q_{t}}(y)=\frac{e^{-s}}{1-e^{-2s}}\bigl(\mathbb{E}_{\nu_{y,s}}[X]-\mathbb{E}_{\mu_{y,s}}[X]\bigr)\,$,
@@
As $\sup_{y\in\mathbb{R}^{d}}{\rm FI}(\nu_{y,t}\,\|\,\mu_{y,t})\leq\operatorname{FI}_{\infty}(\nu\,\|\,\mu)$, our proof is complete. ∎
@@
Proof.  Suppose $\mu$ satisfies the assumptions of Corollary 3.5, i.e., $\mu\propto\exp(-(V+H))$ where $V$ is $\alpha$-strongly convex and $H$ is $L$-Lipschitz. Then it is easy to see that for any $y\in\mathbb{R}^{d}$ and $s\geq 0$ that $\mu_{y,s}$ is strongly log-concave with parameter $\alpha+1/u(s)$ and the log-perturbation $H$ remains unchanged. Thus by [3, Theorem 1.4], $\mu_{y,s}$ satisfies the log-Sobolev inequality with parameter given by (4.4).
@@
Suppose $\mu$ is such that ($\Theta$) holds for some constants $(\theta_{s})_{s\geq 0}$ and that $\mu_{y,t}$ satisfies a log-Sobolev inequality with constant $(\lambda_{t})_{t\geq 0}$ for any $y\in\mathbb{R}^{d}$ and $t\geq 0$. Then $\gamma$-a.s.,
$\|X_{T}-Y_{T}\|\leq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,\eta_{T}\coloneqq\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\int_{0}^{T}d_{T-s}\exp\Bigl(\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,$,
where $d_{T-s}=e^{T-s}\lambda_{u(T-s)}/u(T-s)$ and $u(T-s)\coloneqq e^{2(T-s)}-1$. Assuming that $\eta_{\infty}\coloneqq\lim_{T\to\infty}\eta_{T}$ exists, we have that
$\|T^{\mu}_{\rm KM}-T^{\nu}_{\rm KM}\|_{L^{\infty}(\gamma)}\leq\eta_{\infty}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,.$
We briefly mention that, to the best of our knowledge, Theorem 4.3 establishes a new transport-information inequality of the form
@@
where we recall that $W_{\infty}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}{\rm{ess\,sup}}_{(X,Y)\sim\pi}\|X-Y\|$, where $\Pi(\mu,\nu)$ is the set of joint measures with first- and second-marginal given by $\mu$ and $\nu$ respectively. In particular, we can revisit our previous examples (log-Lipschitz perturbations and strongly log-concave outside a ball) in the context of Theorem 4.3, where we obtain the same constants as Section 3.
@@
Suppose $\mu\propto\exp(-(V+H))$ where $V$ is $\alpha$-strongly convex and $H$ is $L$-Lipschitz. Then Theorem 4.3 holds with constant
@@
Proof.  By Theorem 4.3, we can compute $d_{s}$ and use the bound $\alpha+1/u(s)\geq\alpha$ for all $s\geq 0$ to obtain
@@
$\eta_{T}\leq\exp\Bigl(\frac{L^{2}}{\alpha}+\frac{4L}{\sqrt{\alpha}}\Bigr)\int_{0}^{T}\frac{e^{(T-s)}}{\alpha\,(e^{2(T-s)}-1)+1}\exp\Bigl(\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$
@@
$\eta_{T}\leq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\int_{0}^{T}\frac{e^{2(T-s)}}{(\alpha e^{2(T-s)}+(1-\alpha))^{3/2}}\,\mathrm{d}s\,.$
@@
Proof.  The computations follow verbatim the arguments from Corollary 4.4, using the second part of Theorem 4.3, and are thus omitted. ∎
5 Conclusion
This note establishes, to our knowledge, the first stability properties of the Kim–Milman flow map. Some settings in our work (i.e., log-Lipschitz perturbations of strongly log-concave measures, or having asymptotically convex profiles) have not been analyzed for optimal transport maps or entropic transport maps. We show that our proof technique can be modified to give $L^{\infty}$-type stability bounds as well. In all cases, it is natural to extend these stability properties to measures which live on Riemannian manifolds, or to other flow maps (e.g., the flow map induced by general stochastic interpolants [1, 18] or the Schrödinger bridge flow map [15, 22, 25]).
Acknowledgments
We thank Zhou Fan and Daniel Lacker for helpful discussions. AAP thanks the Foundations of Data Science at Yale University for support. MSZ is funded by an NSERC CGS-D award.

### v2

@@ 3.1.2 Distributions with asymptotically positive convex profiles
As a final example, we turn to a family of distributions whose origins date back to [12, 24]. To define said family, we require the following definitions. For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we define the integrated convexity profile of $f$, denoted $\kappa_{f}:\mathbb{R}_{+}\to\mathbb{R}\cup\{-\infty\}$, to be
$\kappa_{f}(r)\coloneqq\inf\Bigl\{\frac{\langle\nabla{f}(x)-\nabla{f}(y),x-y\rangle}{\|x-y\|^{2}}\ :\ \|x-y\|=r\Bigr\}\,.$
We motivate this definition with the following example.
@@
If $R_{V}=0$, then $\kappa_{V}(r)\geq\alpha_{V}>0$ is equivalent to strong convexity over all of $\mathbb{R}^{d}$. Otherwise for $R_{V}>0$, the potential has integrated convexity profile which might be negative inside $B(0,R_{V})$, while remaining strongly convex outside the ball.
@@
    the right-derivative of g{g} at the origin exists, denoted $g^{\prime}(0)$}.
This function class leads to our final corollary of this section.
@@
Let ${\alpha}>-1$ and $\hat{g}\in\mathcal{G}$, and $\mu\propto\gamma\exp(-h)$ where the integrated convexity profile of $h$ satisfies $\kappa_{h}(r)\geq\alpha-r^{-1}\hat{g}(r)$. Then,
$\Lambda_{\infty}=\frac{1}{1+\alpha}\exp\Bigl(\frac{3\hat{g}^{\prime}(0)}{2\,(1+\alpha)}\Bigr)\,.$
@@
${\Lambda_{T}}$ $\leq\exp\Bigl(-\frac{3\hat{g}^{\prime}(0)}{2\alpha}\bigl((1+\alpha)^{-1}-1)\bigr)\Bigr)\int_{\exp(-T)}^{1}(1+\alpha-\alpha r^{2})^{-3/2}\,\mathrm{d}r\,$,
@@
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq{\alpha_{V}}^{-1}\exp\bigl(O\bigl((L_{V}/{\alpha_{V}})\,(1\vee L_{V}R_{V}^{2})\bigr)\bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
@@ 4 Extension to stronger metrics
To start, we follow the start of the proof of Theorem 3.1 assuming ($\Theta$). It is easy to see that $\gamma$-almost surely, it holds that
@@
We see that the rest of the argument is carried out as before, replacing $\Lambda_{T}$ with
${\eta_{T}\coloneqq}\int_{0}^{T}d_{s}\exp\Bigl(\int_{0}^{{s}}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$ (4.3)
@@
To this end, for a measure $\mu$, $y\in\mathbb{R}^{d}$ and $t\geq 0$, we write
@@
${\rm KL}(\nu\,\|\,\mu)\coloneqq\int{\log\Bigl(\frac{\nu}{\mu}\Bigr)\,\mathrm{d}\nu}\leq\frac{\lambda}{2}\,{\rm FI}(\nu\,\|\,\mu)\,.$
@@
$\nabla\log\frac{\nu Q_{s}}{\mu Q_{s}}(y)=\frac{e^{-s}}{1-e^{-2s}}\bigl(\mathbb{E}_{\nu_{y,s}}[X]-\mathbb{E}_{\mu_{y,s}}[X]\bigr)\,$,
@@
As $\sup_{y\in\mathbb{R}^{d}}{\rm FI}(\nu_{y,s}\,\|\,\mu_{y,s})\leq\operatorname{FI}_{\infty}(\nu\,\|\,\mu)$, our proof is complete. ∎
@@
Proof.  Suppose $\mu$ satisfies the assumptions of Corollary 3.3, i.e., $\mu\propto\exp(-{V+H})$ where $V$ is $\alpha$-strongly convex and $H$ is $L$-Lipschitz. Then it is easy to see that for any $y\in\mathbb{R}^{d}$ and $s\geq 0$ that $\mu_{y,s}$ is strongly log-concave with parameter $\alpha+1/u(s)$ and the log-perturbation $H$ remains unchanged. Thus by [3, Theorem 1.4], $\mu_{y,s}$ satisfies the log-Sobolev inequality with parameter given by (4.4).
@@
Suppose $\mu$ is such that ($\Theta$) holds for some constants $(\theta_{s})_{s\geq 0}$ and that $\mu_{y,t}$ satisfies a log-Sobolev inequality with constant $\lambda_{t}$} for any $y\in\mathbb{R}^{d}$ and $t\geq 0$. Then $\gamma$-a.s.,
$\|T^{\mu}_{\rm KM}-T^{\nu}_{\rm KM}\|_{L^{\infty}(\gamma)}{\leq\eta_{\infty}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\coloneqq\Bigl[\int_{0}^{\infty}\frac{e^{T}\lambda_{T}}{e^{2T}-1}\exp\Bigl(\int_{0}^{T}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}T\Bigr]}\sqrt{\operatorname{FI}_{\infty}(\nu\,\|\,\mu)}\,.$
Remark 4.4.
To the best of our knowledge, Theorem 4.3 establishes a new transport-information inequality of the form
@@
where we recall that $W_{\infty}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}{\rm{ess\,sup}}_{(X,Y)\sim\pi}\|X-Y\|$, where $\Pi(\mu,\nu)$ is the set of joint measures with first- and second-marginal given by $\mu$ and $\nu$ respectively.
We will now revisit our previous examples (log-Lipschitz perturbations and strongly log-concave outside a ball) in the context of Theorem 4.3, where we obtain the same constants as in Section 3.
@@
Suppose $\mu\propto\exp(-{V+H})$ where $V$ is $\alpha$-strongly convex and $H$ is $L$-Lipschitz. Then Theorem 4.3 holds with constant
@@
Proof.  By Lemma 4.1 and Proposition 4.2, we can compute $d_{s}$ and use the bound $\alpha+1/u(s)\geq\alpha$ for all $s\geq 0$ to obtain
@@
$\eta_{T}\leq\exp\Bigl(\frac{L^{2}}{\alpha}+\frac{4L}{\sqrt{\alpha}}\Bigr)\int_{0}^{T}\frac{e^{{s}}}{\alpha\,(e^{2{s}}-1)+1}\exp\Bigl(\int_{0}^{{s}}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$
@@
$\eta_{T}\leq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\int_{0}^{T}\frac{e^{2{s}}}{(\alpha e^{2{s}}+(1-\alpha))^{3/2}}\,\mathrm{d}s\,.$
@@
Proof.  The computations follow verbatim the arguments from Corollary 4.5, using the second part of Proposition 4.2, and are thus omitted. ∎